\begin{thebibliography}{BP{\O}22b}

\bibitem[AGV72]{SGA4}
Michael {Artin}, Alexander {Grothendieck}, and J.~L. {Verdier}.
\newblock {\em {S\'eminaire de g\'eom\'etrie alg\'ebrique du Bois-Marie 1963--1964. Th\'eorie des topos et cohomologie \'etale des sch\'emas. (SGA 4). Un s\'eminaire dirig\'e par M. Artin, A. Grothendieck, J. L. Verdier. Avec la collaboration de N. Bourbaki, P. Deligne, B. Saint-Donat. Tome 1: Th\'eorie des topos. Expos\'es I \`a IV. 2e \'ed.}}, volume 269 of {\em {Lecture Notes in Mathematics}}.
\newblock Springer, Cham, 1972.

\bibitem[AHI24]{AHI-Atiyah}
Toni Annala, Marc Hoyois, and Ryomei Iwasa.
\newblock Atiyah duality for motivic spectra.
\newblock arxiv preprint: \url{https://arxiv.org/abs/2403.01561}, 2024.

\bibitem[AI23]{AnnalaIwasaUnivers}
Toni Annala and Ryomei Iwasa.
\newblock Motivic spectra and universality of {$K$}-theory.
\newblock Arxiv preprint: \url{https://arxiv.org/abs/2204.03434}, 2023.

\bibitem[Ayo07]{Ayoub-thesis}
Joseph Ayoub.
\newblock {\em Les six op{\'e}rations de {Grothendieck} et le formalisme des cycles {\'e}vanescents dans le monde motivique. {I}}, volume 314 of {\em Ast{\'e}risque}.
\newblock Paris: Soci{\'e}t{\'e} Math{\'e}matique de France, 2007.

\bibitem[BLMP]{BLMP}
Federico Binda, Tommy Lundemo, Alberto Merici, and Doosung Park.
\newblock Logarithmic prismatic cohomology, motivic sheaves, and comparison theorems.

\bibitem[BP{\O}22a]{BPOCras}
Federico Binda, Doosung Park, and Paul~Arne {\O}stv{\ae}r.
\newblock Motives and homotopy theory in logarithmic geometry.
\newblock {\em C. R., Math., Acad. Sci. Paris}, 360:717--727, 2022.

\bibitem[BP{\O}22b]{BPO}
Federico Binda, Doosung Park, and Paul~Arne {\O}stv{\ae}r.
\newblock {\em Triangulated categories of logarithmic motives over a field}, volume 433 of {\em Ast{\'e}risque}.
\newblock Paris: Soci{\'e}t{\'e} Math{\'e}matique de France (SMF), 2022.

\bibitem[BP{\O}23]{BPO-SH}
Federico Binda, Doosung Park, and Paul~Arne {\O}stv{\ae}r.
\newblock Logarithmic motivic homotopy theory.
\newblock ArXiv preprint: \url{https://arxiv.org/abs/2303.02729}, 2023.

\bibitem[CD09]{CiDeg-LocalStableHomAlg}
Denis-Charles Cisinski and Fr{\'e}d{\'e}ric D{\'e}glise.
\newblock Local and stable homological algebra in {G}rothendieck abelian categories.
\newblock {\em Homology Homotopy Appl.}, 11(1):219--260, 2009.

\bibitem[CD16]{CD-etale}
Denis-Charles Cisinski and Frédéric Déglise.
\newblock \'etale motives.
\newblock {\em Compositio Math.}, 152(3):556–666, 2016.

\bibitem[Eke85]{EkedahlII}
Torsten Ekedahl.
\newblock On the multiplicative properties of the de {R}ham—{W}itt complex. ii.
\newblock {\em Ark. Mat.}, 23:53--102, 1985.

\bibitem[ESS21]{ertl2021integral}
Veronika Ertl, Atsushi Shiho, and Johannes Sprang.
\newblock Integral {\it p}-adic cohomology theories for open and singular varieties, 2021.
\newblock ArXiv preprint \url{https://arxiv.org/abs/2105.11009}.

\bibitem[GL00]{GL}
Thomas Geisser and Marc Levine.
\newblock The {$K$}-theory of fields in characteristic {$p$}.
\newblock {\em Invent. Math.}, 139:459--493, 2000.

\bibitem[Gro85]{Gros1985}
Michel Gros.
\newblock Classes de {C}hern et classes de cycles en cohomologie de {H}odge-{W}itt logarithmique.
\newblock {\em M{\'e}m. Soc. Math. Fr.}, 21:1--87, 1985.

\bibitem[Hir64]{Hironaka}
Heisuke Hironaka.
\newblock Resolution of singularities of an algebraic variety over a field of characteristic zero: I.
\newblock {\em Ann. of Math.}, 79(1):109--203, 1964.

\bibitem[Hov01]{HoveySpectra}
Mark Hovey.
\newblock Spectra and symmetric spectra in general model categories.
\newblock {\em J. Pure Appl. Algebra}, 165(1):63--127, 2001.

\bibitem[HS21]{HS2021}
Katharina H\"ubner and Alexander Schmidt.
\newblock The tame site of a scheme.
\newblock {\em Inv. Math.}, 223:397--443, 2021.

\bibitem[H{\"u}b21]{Huebner2021}
Katharina H{\"u}bner.
\newblock The adic tame site.
\newblock {\em Doc. Math.}, 26:873--945, 2021.

\bibitem[Kou24]{amine}
Amine Koubaa.
\newblock The {C}artier operator on differentials of discretely ringed adic spaces and purity in the tame cohomology, 2024.
\newblock ArXiv preprint: \url{https://arxiv.org/abs/2408.02542}.

\bibitem[Lor09]{lorenzon}
Pierre Lorenzon.
\newblock Logarithmic hodge–witt forms and hyodo–kato cohomology.
\newblock {\em Journal of Algebra}, 249:247--265, 2009.

\bibitem[Mat17]{matsuue}
Hironori Matsuue.
\newblock On relative and overconvergent de rham–witt cohomology for log schemes.
\newblock {\em Mathematische Zeitschrift}, 286:19--87, 2017.

\bibitem[Mer22]{mericicrys}
Alberto Merici.
\newblock A motivic integral $p$-adic cohomology.
\newblock Arxiv preprint: \url{{https://arxiv.org/abs/2211.14303}}, 2022.

\bibitem[Mok93]{mokrane}
Abdellah Mokrane.
\newblock La suite spectrale des poids en cohomologie de {H}yodo-{K}ato.
\newblock {\em Duke Mathematical Journal}, 72:301--337, 1993.

\bibitem[MV99]{MV}
Fabien Morel and Vladimir Voevodsky.
\newblock {${\bf A}^1$}-homotopy theory of schemes.
\newblock {\em Inst. Hautes \'Etudes Sci. Publ. Math.}, 90:45--143 (2001), 1999.

\bibitem[MVW06]{MVW}
Carlo Mazza, Vladimir Voevodsky, and Charles Weibel.
\newblock {\em Lecture notes on motivic cohomology}, volume~2 of {\em Clay Math. Monographs}.
\newblock American Mathematical Society, 2006.

\bibitem[Ogu18]{ogu}
Arthur Ogus.
\newblock {\em Lectures on logarithmic algebraic geometry}, volume 178 of {\em Cambridge Stud. Adv. Math.}
\newblock Cambridge Univ. Press, 2018.

\bibitem[Par23]{DoosungA1invlog}
Doosung Park.
\newblock On the log motivic stable homotopy groups.
\newblock to appear in Homology, Homotopy and Applications, 2023.

\bibitem[Rob15]{robalo}
Marco Robalo.
\newblock K-theory and the bridge from motives to noncommutative motives.
\newblock {\em Adv. Math.}, 269:399--550, 2015.

\bibitem[Sai23]{shujilog}
Shuji Saito.
\newblock Reciprocity sheaves and logarithmic motives.
\newblock {\em Compositio Math.}, 159:355--379, 2023.

\bibitem[{Sta}16]{stacks-project}
{Stacks Project Authors}.
\newblock {\itshape {S}tacks {P}roject}.
\newblock \url{http://stacks.math.columbia.edu}, 2016.

\bibitem[Voe00]{VoevTriangCat}
Vladimir Voevodsky.
\newblock Triangulated categories of motives over a field.
\newblock In {\em Cycles, transfers, and motivic homology theories}, volume 143 of {\em Ann. of Math. Stud.}, pages 188--238. 2000.

\end{thebibliography}

\documentclass[11pt, a4paper]{amsart}

\usepackage{amscd,verbatim}
\usepackage{amssymb}

\usepackage{bm}

\usepackage{amssymb, amsmath}
\usepackage[all]{xy}
\usepackage[inline]{enumitem}
\usepackage{hyphenat}
\usepackage[colorlinks,linkcolor=blue,citecolor=blue,urlcolor=red]{hyperref}
\usepackage[dvipsnames]{xcolor}
\usepackage{mathtools}
\usepackage{tikz-cd}
\usetikzlibrary{cd,decorations.pathreplacing}
\usepackage[labelformat=empty]{caption}
\usepackage{xfrac}
 \usepackage[ left=3cm, right=3cm, top = 2.8cm, bottom = 2.8cm ]{geometry}
\usepackage{stmaryrd}
\usepackage{bbm}

\makeatletter
\def\@tocline#1#2#3#4#5#6#7{\relax
  \ifnum #1>\c@tocdepth 
  \else
    \par \addpenalty\@secpenalty\addvspace{#2}%
    \begingroup \hyphenpenalty\@M
    \@ifempty{#4}{%
      \@tempdima\csname r@tocindent\number#1\endcsname\relax
    }{%
      \@tempdima#4\relax
    }%
    \parindent\z@ \leftskip#3\relax \advance\leftskip\@tempdima\relax
    \rightskip\@pnumwidth plus4em \parfillskip-\@pnumwidth
    #5\leavevmode\hskip-\@tempdima
      \ifcase #1
      \or\or \hskip 2em \or \hskip 2em \else \hskip 3em \fi%
      #6\nobreak\relax
    \dotfill\hbox to\@pnumwidth{\@tocpagenum{#7}}\par
    \nobreak
    \endgroup
  \fi}
\makeatother

\newcommand{\A}{\mathbf{A}}

\newcommand{\G}{\mathbf{G}}

\renewcommand{\P}{\mathbf{P}}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\F}{\mathbb{F}}

\newcommand{\sP}{\mathcal{P}}

\newcommand{\Cor}{\operatorname{\mathbf{Cor}}}

\newcommand{\Map}{\operatorname{map}}

\newcommand{\uHom}{\operatorname{\underline{Hom}}}

\newcommand{\Spec}{\operatorname{Spec}}

\newcommand{\pro}[1]{\text{\rm pro}_{#1}\text{\rm--}}
\newcommand{\tr}{{\operatorname{tr}}}

\newcommand{\eff}{{\operatorname{eff}}}

\newcommand{\Nis}{{\operatorname{Nis}}}
\newcommand{\et}{{\operatorname{\acute{e}t}}}

\newcommand{\Sym}{{\operatorname{Sym}}}

\renewcommand{\lim}{\operatornamewithlimits{\varprojlim}}
\newcommand{\colim}{\operatornamewithlimits{\varinjlim}}

\newcommand{\ol}{\overline}

\renewcommand{\phi}{\varphi}
\renewcommand{\epsilon}{\varepsilon}

\newcommand{\CI}{\operatorname{\mathbf{CI}}}

\newcommand{\bcube}{{\ol{\square}}}

\newcommand{\M}{\mathbf{M}}

\newcounter{spec}
{\end{list}}%

\newtheorem{lemma}{Lemma}[section]
\newtheorem{thm}[lemma]{Theorem}

\newtheorem{prop}[lemma]{Proposition}

\newtheorem{cor}[lemma]{Corollary}

\theoremstyle{definition}

\theoremstyle{remark}

\newtheorem{rmk}[lemma]{Remark}

\newtheorem{example}[lemma]{Example}

\newtheorem{nota}[lemma]{Notation}
\numberwithin{equation}{section}

\numberwithin{equation}{lemma}

\colorlet{LightRubineRed}{RubineRed!70!}

\usepackage[color = blue!20, bordercolor = black, textsize = tiny]{todonotes}

\usepackage{relsize} 
\usepackage[bbgreekl]{mathbbol} 
\usepackage{amsfonts} 
\DeclareSymbolFontAlphabet{\mathbb}{AMSb} 
\DeclareSymbolFontAlphabet{\mathbbl}{bbold} 

\DeclareSymbolFontAlphabet{\mathbbl}{bbold}

\def\lSm{\mathbf{lSm}}
\def\SmlSm{\mathbf{SmlSm}}
\def\Sm{\mathbf{Sm}}

\newcounter{elno}

\begin{document}

\def\THH{\operatorname{THH}}
\def\TC{\operatorname{TC}}
\def\TCmin{\operatorname{TC}^-}
\def\TP{\operatorname{TP}}
\def\HH{\operatorname{HH}}
\def\HC{\operatorname{HC}}
\def\HCmin{\operatorname{HC}^-}
\def\HP{\operatorname{HP}}

\def\Fil{\operatorname{Fil}}
\def\Gr{\operatorname{Gr}}
\def\gr{\operatorname{gr}}

\def\QSyn{\operatorname{QSyn}}
\def\QRSPerfd{\operatorname{QRSPerfd}}
\def\lQSyn{\operatorname{lQSyn}}
\def\lQRSPerfd{\operatorname{lQRSPerfd}}

\def\syn{\mathrm{syn}}
\def\Fsyn{\mathrm{Fsyn}}
\def\Fet{\mathrm{F\acute{e}t}}

\def\LogRec{\operatorname{\mathbf{LogRec}}}
\def\Ch{\operatorname{\mathrm{Ch}}}

\def\ltr{\mathrm{ltr}}

\def\kX{\mathfrak{X}}
\def\kY{\mathfrak{Y}}

\def\otCIsp{\otimes_{\CI}^{sp}}
\def\otCINissp{\otimes_{\CI}^{\Nis,sp}}

\def\tL{\tilde{L}}
\def\tX{\tilde{X}}
\def\tY{\tilde{Y}}
\def\tF{\widetilde{F}}
\def\tG{\widetilde{G}}

\def\Sh{\operatorname{\mathbf{Shv}}}
\def\PSh{\operatorname{\mathbf{PSh}}}
\def\Shltr{\operatorname{\mathbf{Shv}_{dNis}^{ltr}}}
\def\Shlog{\operatorname{\mathbf{Shv}_{dNis}^{log}}}
\def\Shvlog{\operatorname{\mathbf{Shv}^{log}}}
\def\Sm{\operatorname{\mathrm{Sm}}}
\def\SmlSm{\operatorname{\mathrm{SmlSm}}}
\def\lSm{\operatorname{\mathrm{lSm}}}
\def\FlSm{\operatorname{\mathrm{FlSm}}}
\def\FlQSm{\operatorname{\mathrm{FlQSm}}}
\def\FlQSyn{\operatorname{\mathrm{FlQSyn}}}
\def\lCor{\operatorname{\mathrm{lCor}}}
\def\SmlCor{\operatorname{\mathrm{SmlCor}}}
\def\PShltr{\operatorname{\mathbf{PSh}^{ltr}}}
\def\PShlog{\operatorname{\mathbf{PSh}^{log}}}
\def\logCI{\mathbf{logCI}} 

\def\Mod{\operatorname{Mod}}
\newcommand{\DM}[1][]{\operatorname{\mathcal{DM}}_{#1}}
\newcommand{\DMeff}[1][]{\operatorname{\mathcal{DM}}^{\eff}_{#1}}
\newcommand{\DA}[1][]{\operatorname{\mathcal{DA}}_{#1}}
\newcommand{\DAeff}[1][]{\operatorname{\mathcal{DA}}^{\eff}_{#1}}
\newcommand{\FDA}[1][]{\operatorname{\mathcal{FDA}_{#1}}}
\newcommand{\FDAeff}[1][]{\operatorname{\mathcal{FDA}^{\eff}_{#1}}}
\newcommand{\SH}[1][]{\operatorname{\mathcal{SH}_{#1}}}
\newcommand{\logSH}[1][]{\operatorname{\mathbf{log}\mathcal{SH}_{#1}}}
\newcommand{\logDA}[1][]{\operatorname{\mathbf{log}\mathcal{DA}_{#1}}}
\newcommand{\logDAeff}[1][]{\operatorname{\mathbf{log}\mathcal{DA}^{\eff}_{#1}}}
\newcommand{\logDM}[1][]{\operatorname{\mathbf{log}\mathcal{DM}_{#1}}}
\newcommand{\logDMeff}[1][]{\operatorname{\mathbf{log}\mathcal{DM}^{\eff}_{#1}}}
\newcommand{\logFDA}[1][]{\operatorname{\mathbf{log}\mathcal{FDA}_{#1}}}
\newcommand{\logFDAeff}[1][]{\operatorname{\mathbf{log}\mathcal{FDA}^{\eff}_{#1}}}
\newcommand{\WOmega}{\operatorname{\mathcal{W}\Omega}}
\def\Log{\operatorname{\mathcal{L}\textit{og}}}
\def\Rsc{\operatorname{\mathcal{R}\textit{sc}}}
\def\Pro{\mathrm{Pro}\textrm{-}}
\def\pro{\mathrm{pro}\textrm{-}}
\def\dg{\mathrm{dg}}
\def\plim{\mathrm{``lim"}}
\def\ker{\mathrm{ker}}
\def\coker{\mathrm{coker}}
\def\PrL{\mathcal{P}\mathrm{r^L}}
\def\PrLo{\mathcal{P}\mathrm{r^{L,\otimes}}}
\def\Spt{\mathcal{S}\mathrm{pt}}
\def\PSpt{\mathrm{Pre}\mathcal{S}\mathrm{pt}}
\def\Fun{\mathrm{Fun}}
\def\Sym{\mathrm{Sym}}
\def\CAlg{\mathrm{CAlg}}
\def\Poly{\mathrm{Poly}}
\def\Cat{\mathrm{Cat}}

\def\Alb{\operatorname{Alb}}
\def\bAlb{\mathbf{Alb}}
\def\Gal{\operatorname{Gal}}

\def\hofib{\mathrm{hofib}}
\def\fib{\mathrm{fib}}
\def\triv{\mathrm{triv}}
\def\ABl{\mathcal{A}\textit{Bl}}
\def\divsm#1{{#1_\mathrm{div}^{\mathrm{Sm}}}}

\def\cA{\mathcal{A}}
\def\cB{\mathcal{B}}
\def\cC{\mathcal{C}}
\def\cD{\mathcal{D}}
\def\cE{\mathcal{E}}
\def\cF{\mathcal{F}}
\def\cG{\mathcal{G}}
\def\cH{\mathcal{H}}
\def\cI{\mathcal{I}}
\def\cS{\mathcal{S}}
\def\cM{\mathcal{M}}
\def\cO{\mathcal{O}}
\def\cP{\mathcal{P}}

\def\tcA{\widetilde{\mathcal{A}}}
\def\tcB{\widetilde{\mathcal{B}}}
\def\tcC{\widetilde{\mathcal{C}}}
\def\tcD{\widetilde{\mathcal{D}}}
\def\tcE{\widetilde{\mathcal{E}}}
\def\tcF{\widetilde{\mathcal{F}}}

\def\one{\mathbbm{1}}

\def\XP{X \backslash \sP}
\def\M0a{{}^t\cM_0^a}
\newcommand{\Ind}{{\operatorname{Ind}}}

\def\Xkbar{\overline{X}_{\overline{k}}}
\def\dx{{\rm d}x}

\newcommand{\dNis}{{\operatorname{dNis}}}
\newcommand{\loget}{{\operatorname{l\acute{e}t}}}
\newcommand{\ket}{{\operatorname{k\acute{e}t}}}
\newcommand{\ABNis}{{\operatorname{AB-Nis}}}
\newcommand{\sNis}{{\operatorname{sNis}}}
\newcommand{\sZar}{{\operatorname{sZar}}}
\newcommand{\set}{{\operatorname{s\acute{e}t}}}
\newcommand{\cofib}{\mathrm{Cofib}}

\newcommand{\Gmlog}{\G_m^{\log}}
\newcommand{\Gmlogred}{\overline{\G_m^{\log}}}

\newcommand{\varcolim}{\mathop{\mathrm{colim}}}
\newcommand{\varlim}{\mathop{\mathrm{lim}}}
\newcommand{\tensor}{\otimes}

\newcommand{\eq}[2]{\begin{equation}\label{#1}#2 \end{equation}}
\newcommand{\eqalign}[2]{\begin{equation}\label{#1}\begin{aligned}#2 \end{aligned}\end{equation}}

\def\varplim#1{\text{``}\varlim_{#1}\text{''}}
\def\det{\mathrm{d\acute{e}t}}
\def\federem#1{\begin{color}{teal}{#1}\end{color}}
\def\tomrem#1{\begin{color}{purple}{#1}\end{color}}

\author{Alberto Merici}
\address{Institut f\"ur Mathematik, Universit\"at Heidelberg\\ MATHEMATIKON, Im Neuenheimer Feld 205, 69120  Heidelberg, Germany.}
\email[A. Merici]{merici@mathi.uni-heidelberg.de}

\thanks{A.M. is supported by Horizon Europe's Marie
Sk{\l}odowska-Curie Action PF 101103309 ``MIPAC''\\
MSC classes: 14F30 (Primary), 14F42, 19E15}
\title[Motivic p-adic tame cohomology]{Motivic $p$-adic tame cohomology}

\begin{abstract} 
We construct a comparison functor between ($\A^1$-local) tame motives and ($\bcube$-local) log-\'etale motives over a field $k$ of positive characteristic. This generalizes Binda--Park--{\O}stv{\ae}r's comparison for the Nisnevich topology. As a consequence, we construct an $E_\infty$-ring spectrum $H\Z/p^m$ representing mod $p^m$ tame motivic cohomology: the existence of this ring spectrum and the usual properties of motives imply some results on tame motivic cohomology and a comparison with log \'etale motivic cohomology.
\end{abstract}
\maketitle
\section{Introduction}
Let $k$ be a field of characteristic $p$ and let $\Sm_k$ (resp $\lSm_k$) denote the category of (log) smooth (log) schemes over $k$, and $\bcube$ the log scheme $(\P^1,\infty)$, which ideally sits in-between $\A^1$ and $\P^1$ and represents ``sections on $\A^1$ that behave well at $\infty$'' (see \cite[Figure 1]{BPOCras}). The triangulated category of effective logarithmic motives over a field $\logDMeff(k)$ and its non-effective counterpart $\logDM(k)$  were introduced in \cite{BPO} generalizing the $\A^1$-invariant category of Voevodsky motives of \cite{VoevTriangCat} (see also \cite{MVW}), by choosing a suitable version of the Nisnevich topology for log schemes (the dividing Nisnevich topology, $\dNis$) and localizing at $\bcube$, with the scope of studying non-$\A^1$-invariant cohomology theories of log schemes. Later, for $S$ any quasi-compact quasi-separated log scheme, the $\bcube$-homotopy category $\mathbf{log}\mathcal{H}(S)$ and the $S^1$ stable and $\P^1$-stable counterparts $\logSH^{S^1}(S)$ and $\logSH(S)$ were introduced by the same authors in \cite{BPO-SH}, generalizing Morel---Voevodsky $\A^1$-homotopy categories $\mathcal{H}(S)$, $\SH^{S^1}(S)$ and $\SH(S)$ of \cite{MV} (see also \cite{Ayoub-thesis}), in order to study cohomology theories of log schemes represented by sheaves of spectra. In \cite[Theorem 8.2.11]{BPO} and \cite[Theorem 4.4]{DoosungA1invlog}, it was shown that the functor $\omega\colon \lSm_k\to \Sm_k$ that sends a log scheme $(X,\partial X)$ to $X-|\partial X|$ induces fully faithful functors:\[
\DMeff(k)\xrightarrow{\omega^*} \logDMeff(k)\quad \SH^{S^1}(k)\xrightarrow{\omega^*} \logSH^{S^1}(k),
\]
both characterized by the fact that for all $G\in \logDMeff(k)$ (or $\logSH^{S^1}(k)$) and all $(\ol{X},\partial X)\in \lSm_k$\[
(\omega^*G)(X,\partial X) = G(X-|\partial X|).
\]
This justifies the choice of one notation to denote both functors: it will be clear from the context which one is considered.
The functors $\omega^*$ have right adjoints $\omega_*$ by design (the $\A^1$-colocalization). If $k$ satisfies resolutions of singularities as in Notation \ref{nota:RS} (analogous to \cite[Main Theorem I and II]{Hironaka}), then these functors send the motive of a smooth scheme $X$ to the log motive of any smooth log compactification $(\ol{X},\partial X)$ as in Remark \ref{rmk:comp}, so the functors $\omega_*$ are both characterized by the fact that for all $F\in \logDMeff(k)$ (or $\logSH^{S^1}(k)$) and all $X\in \Sm_k$ with smooth log compactification $(\ol{X},\partial X)$\[
(\omega_*)F(X) = F(\ol{X},\partial X).
\]
In \cite[Remark 6.3]{mericicrys}, it was shown that this adjunction cannot be promoted to an adjunction between \'etale and log \'etale motives: in fact, if $X\in \Sm_k$ and $Y\to X$ is an \'etale cover, then the log \'etale sheafification of the \v Cech nerve $L_{\loget}\omega^*M(Y^\bullet)\to L_{\loget}\omega^*M(X)$ need not be an equivalence: the counterexample (which comes from \cite{ertl2021integral}) is an Artin--Schreier cover, which has wild ramification.

In this article, we show that by substituting the \'etale topology with the tame topology defined by H\"ubner--Schmidt in \cite{HS2021}, we indeed have a positive result, namely:
    \begin{thm}[see Theorem \ref{thm:comparison}]\label{thm:mot-intro}
    Let $k$ be a perfect field that satisfies resolutions of singularities as in Notation \ref{nota:RS}. Then the adjunction\[
    \begin{tikzcd}
    \SH^{S^1}(k) \ar[r,"\omega^*", shift left = 2] &\logSH^{S^1}(k)\ar[l,"\omega_*"]
    \end{tikzcd}
    \]
    induces an adjunction\[
    \begin{tikzcd}
    \SH^{S^1}_t(k) \ar[r,"\omega^*_t", shift left = 2] &\logSH^{S^1}_{\loget}\ar[l,"\omega_*^t"](k)
    \end{tikzcd}
    \]
fitting in commutative diagrams:\[
    \begin{tikzcd}
    \SH^{S^1}(k) \ar[r,"\omega^*"]\ar[d,"L_{(\A^1,t)}"] &\logSH^{S^1}(k)\ar[d,"L_{(\bcube,\loget)}"] && 
    \SH^{S^1}(k) &\logSH^{S^1}(k)\ar[l,"\omega_*"]\\
     \SH^{S^1}_t(k) \ar[r,"\omega^*_t"] &\logSH^{S^1}_{\loget}(k) &&
    \SH^{S^1}_t(k) \ar[u,"i_{(\A^1,t)}"]&\logSH^{S^1}_{\loget}\ar[l,"\omega_*^t"](k)\ar[u,"i_{(\bcube,\loget)}"]
    \end{tikzcd}
\]
and similarly for $\DMeff$.
\end{thm}
This result should be compared to \cite{AHI-Atiyah}, where the authors also consider the $\A^1$-colocalization of the motivic spectra constructed in \cite{AnnalaIwasaUnivers}, and get an inclusion $\SH(k)\to \mathbf{Mod}_{1_{\A^1}}\mathcal{MS}(k)$ that sends the motive of $X$ to the total fiber of the Gysin map induced the inclusion of the boundary $|\partial X|\hookrightarrow \ol{X}$ of any smooth log compactification $(\ol{X},\partial X)$ of $X$ as in Remark \ref{rmk:comp} (see \cite[Remark 6.26]{AHI-Atiyah}). For the same reasons explained in \cite[Remark 6.3]{mericicrys}, this cannot be promoted to the \'etale version of $\mathcal{MS}$, and our result shows a promising first step in promoting this in the tame setting. 
We remark that, on the other hand, the motivic spectra of \cite{AnnalaIwasaUnivers} are intrinsically $\P^1$-stable, while the result of Theorem \ref{thm:mot-intro} holds already in an $S^1$-stable setting.

One main application of the previous theorem is the following: for a fixed $m$, let $\nu_m(i)$ denote the mod $p^m$ motivic sheaves of Bloch--Illusie--Milne (to avoid confusion, we will not refer to them as the logarithmic de Rham--Witt sheaves, as they are not sheaves on logarithmic schemes): they are strictly $\A^1$-invariant Nisnevich sheaves with transfers (so $\nu_m(i)[0]\in \DMeff(k)$), but not strictly $\A^1$-invariant \'etale sheaves. In \cite{mericicrys}, we showed that the cohomology of the Rham--Witt sheaves with log poles $W_m\Lambda^n$ of Hyodo--Kato (see \cite{mokrane} or \cite{matsuue}) are representable in the category $\logDMeff[\loget](k)$, and therefore $L_{\loget}\omega^*(\nu_m(i)[0])\simeq \widetilde{\nu_m(i)}[0]$ (see Example \ref{ex:drW}), where $\widetilde{\nu_m(i)}$ is the log \'etale sheaf
\begin{equation}\label{eq:nu-tilde}
\widetilde{\nu_m(i)}\colon (X,\partial X)\mapsto \nu_m(i)(X-|\partial X|).
\end{equation}
As a consequence, we immediately deduce that the presheaves $H^1_t(-/k,\nu_m(n))$ are $\A^1$-invariant (see Example \ref{ex:drW}). Moreover, we deduce the $\A^1$-invariance and $\P^1$-stability of higher tame motivic cohomology by the purity result of Koubaa \cite[Theorem 1.3.1]{amine}, with the same assumptions (RS1) and (RS2) as in \ref{nota:RS}. Putting everything together we show that
\begin{thm}[see \eqref{eq:A1-inv}, Lemma \ref{lem:pbf} and Remark \ref{rmk:GL}]\label{thm:motivic-dRW} 
Let $k$ be a perfect field of characteristic $p$ that satisfies resolutions of singularities as in \ref{nota:RS}. For all $m$, the object \[\{\nu_m(i)[-i]\}_{i\geq 0}\in \mathrm{GrCAlg}(\cD(\Sh_{t}(\Sm_k,\Z/p^m)))\] builds up to an $E_\infty$-ring spectrum $H\Z/p^m$ in $\DM[t](k,\Z/p^m)$ such that for all $X\in \Sm_k$ we have that\[
\Map(\Sigma^{\infty}(X), \Sigma^{p,q}H\Z/p^m) \simeq R\Gamma_t(X,\nu_m(q))[p-q]
\]
Moreover, $H\Z/p^m$ is the unit of $\DM[t](k,\Z/p^m)$. 
\end{thm}
The choice of the terminology $H\Z/p^m$ is justified by \cite{GL}, in fact \emph{a posteriori} $H\Z/p^m$ is the image of motivic cohomology via the localization $\DM(k,\Z/p^m)\to \DM[t](k,\Z/p^m)$.  Using the motivic properties of $\DM$ (see Remark \ref{rmk:pbf-gysin}), we deduce immediately the result:
\begin{thm}\label{thm:main-intro} Let $k$ be a perfect field of characteristic $p$ that satisfies resolutions of singularities as in \ref{nota:RS}. For all $m\geq 1$ and $q\geq 0$, for $X\in \Sm_k$, $\cE\to X$ a vector bundle of rank $r+1$, then we have an isomorphism
        \[
        H^q_t(\cE/k,\nu_m(n))\cong H^q_t(X/k,\nu_m(n))
        \]
        and if $\P(\cE)$ is the associated projective bundle, the Chern classes induce an isomophism\[
        H^q_t(\P(\cE)/k,\nu_m(n))\cong \bigoplus_{i=0}^r H^{q-i}_t(X/k,\nu_m(n-i)).
        \]
\end{thm}
The case $n=0$ of Theorem \ref{thm:motivic-dRW} has already been proved in \cite[15.4]{HS2021} (still assuming that $k$ satisfies resolutions of singularities), deducing it from the adic versions \cite[Corollary 14.5 and 14.6]{Huebner2021}. 
Finally, another interesting application of our result is the following comparison
\begin{thm}\label{thm:comparison-logetale}
    Let $k$ be a field of characteristic $p$ satisfying resolutions of singularities as in \ref{nota:RS}. For all $X\in \Sm_k$ with smooth log compactification $(\ol{X},\partial X)$ as in Remark \ref{rmk:comp} we have that\[
R\Gamma_t(X,\nu_m(n))\simeq R\Gamma_{\loget}((\ol{X},\partial X),\widetilde{\nu_m(n)}),
\]
where $\widetilde{\nu_m(n)}$ is as in \eqref{eq:nu-tilde}.
\end{thm}
We remark that this result requires a very special property of $\nu_m(n)$, \emph{i.e.} the fact that $L_{\loget}\omega^*(\nu_m(n)[0])\simeq \widetilde{\nu_m(n)}[0]$ is $(\bcube,\loget)$-local: we do not know whether to expect this to hold in general for any $(\A^1,\Nis)$-local object.
\subsection*{Future perspectives} 

We expect Theorem \ref{thm:mot-intro} to hold without the assumption on resolutions of singularities. In fact, it was shown in \cite{BLMP} that log prismatic and syntomic cohomology give rise to motivic spectra in $\logSH_{\loget}(S)$ for any quasisyntomic $p$-adic scheme $S$. We expect a similar result as Theorem \ref{thm:mot-intro} to hold for a general base scheme $S$, so that the object $\omega_*^t\mathbf{E}^{\syn}$ can be studied as a tame version of syntomic cohomology. 

\subsection*{Acknowledgements}
	The author would like to thank F. Binda, D. Park, T. Lundemo, K. H\"ubner, A. Schmidt, S. Saito and P.A. {\O}stv{\ae}r for many valuable discussions and comments on earlier versions of the results in this paper, and Joseph Ayoub for pointing out a gap in a previous version of Section 3. He also thanks A. Koubaa for sharing his version of \cite{amine}. He also thanks the anonymous referee for a meticulous analysis of the paper, providing helpful comments which filled some small gaps in the arguments and led to an improved presentation. This project is supported by the MSCA-PF \emph{MIPAC} carried out at the University of Milan. The author is very thankful for the hospitality and the great work environment.

\section{Tame and log-\'etale motives}

We recall the definition of the tame site of \cite{HS2021}. For $S$ a scheme and $X$ an $S$-scheme, the tame site $(X/S)_t$ was defined as the site whose underlying category is $X_{\et}$, and $\{U_i\to U\}$ is a tame cover if an only if for all $x\in U$ and all $S$-valuations $v$ on $k(x)$, there is $y\in U_i$ for some $i$ lying over $x$ and a valuation $w$ on $k(y)$ extending $v$ such that the extension of valued fields $(k(x),v)\to (k(y),w)$ is tame, \emph{i.e.} the ramification index of $\cO_v\to \cO_w$ is prime to the characteristic of the residue field of $\cO_v$. In the rest of the paper, we will only consider the case $S=\Spec(k)$ where $k$ is a perfect field of characteristic $p$, therefore to ease the notation, for $X$ a $k$-scheme we will write $H^*_t(X,-)$ for $H^*_t(X/k,-)$.

We recall the most important properties of tame cohomology:
\begin{enumerate}
    \item By construction, every tame sheaf is a Nisnevich sheaf
    \item If $X$ is a quasi-compact $k$-scheme every tame cover admits a finite subcover \cite[Theorem 4.1]{HS2021}, therefore it is enough to consider covers $U\to X$ where $U=\sqcup U_i$ for a finite cover $\{U_i\to X\}$.
    \item On quasi-compact quasi-separated $k$-schemes, tame cohomology commutes with filtered colimits of sheaves \cite[Theorem 4.5]{HS2021} and cofiltered limits of schemes with affine transition maps \cite[Theorem 4.7]{HS2021}
    \item If $F$ is an \'etale sheaf of $\Z/m\Z$-modules where $m$ is prime to $p$, then $H^q_t(X,F)=H^q_{\et}(X,F)$ for every $X$ quasi-compact $k$-scheme \cite[Proposition 8.1]{HS2021}.
    \item If $X$ is proper over $k$ with the property that every finite set of points is contained in an affine open, then for every tame sheaf $F$ of abelian groups $H^q_t(X,F) = H^q_{\et}(X,a_{\et}F)$ \cite[Proposition 8.2]{HS2021}.
\end{enumerate}

Let $\Cor_k$ be the category of finite correspondences over $k$. By \cite[Lemma 16.1]{HS2021} the category of tame sheaves with transfers $\Sh_t(\Cor_k)$ is a Grothendieck abelian category and for every $F\in \Sh_t(\Cor_k)$ the tame cohomology presheaves \[
U\mapsto H^q_t(U,F)
\]
are presheaves with transfers. Therefore, similarly to Voevodsky's construction, one gets the stable $\infty$-category of effective tame motives and of tame motivic sheaves of $S^1$-spectra as\[
\DMeff[t](k):= L_{(\A^1,t)}\cD(\Sh_t(\Cor_k))\quad \SH_t^{S^1}(k):= L_{(\A^1,t)}\Sh_t(\Sm_k,\Spt).
\]
As usual, the graph functor $\Sm_k\hookrightarrow \Cor_k$ and the Dold-Kan functor $\cD(\Z)\to \Spt$ induce a localization\[
\SH^{S^1}_t(k)\to \DMeff[t](k)
\] 
and there are evident localizations\[
\begin{tikzcd}
\DMeff[\Nis](k)\ar[r,"L_{(\A^1,t)}"]&
\DMeff[t](k)\ar[r,"L_{(\A^1,\et)}"]&\DMeff[\et](k)\\
\SH^{S^1}_{\Nis}(k)\ar[r,"L_{(\A^1,t)}"]&
\SH^{S^1}_t(k)\ar[r,"L_{(\A^1,\et)}"]&\SH^{S^1}_{\et}(k)\\
\end{tikzcd}
\]
with right adjoints\[
\begin{tikzcd}
\DMeff[\et](k)\ar[r,"i_{(\A^1,\et)}"]&
\DMeff[t](k)\ar[r,"i_{(\A^1,t)}"]&\DMeff[\Nis](k)\\
\SH^{S^1}_{\et}(k)\ar[r,"i_{(\A^1,\et)}"]&
\SH^{S^1}_t(k)\ar[r,"i_{(\A^1,t)}"]&\SH^{S^1}_{\Nis}(k)
\end{tikzcd}
\]
For $X\in \Sm_k$ and $\tau$ any of the topologies above, we let $\cM^\tau(X)\in \DMeff[\tau](k)$ and $\cH^\tau(X)\in \SH^{S^1}_{\tau}(k)$ the image of the Yoneda funtcor. By sheafifying \cite[Theorem 3.2.23]{MV} we get a fiber sequence\[
\cH^t(X-Z)\to \cH^t(X)\to \cH^t(\mathrm{Th}_{N_Z}).
\]
\begin{rmk}\label{rmk:pbf-gysin}
    In the case with transfers, by \cite[Properties 14.5.]{MVW} we deduce that
\begin{enumerate}
    \item \label{item:1} If $\cE \to X$ is a vector bundle of rank $r+1$, we have an equivalence\[
        \cM^t(\cE)\xrightarrow{\simeq} \cM^t(X),
    \]
    and if $\P(\cE)$ is the associated projective bundle, the Chern classes induce an equivalence\[
    \bigoplus_{i=0}^r \cM^t(X)\otimes \cM^t(\P^1,i_0)\simeq \cM^t(\P(\cE))
    \]
    where $\cM^t(\P^1,i_0)$ is the complement of the splitting of $\cM^t(\P^1)\to \cM^t(k)$ given by the zero section.
    \item \label{item:2} Let $X$ be a smooth scheme over $k$ and $Z$ a smooth closed
subscheme of $X$ of codimension $c$. Then there is a Gysin fiber sequence\[
\cM^t(X-Z)\to \cM^t(X)\to \cM^t(Z)\otimes \cM^t(\P^1,i_0)^{\otimes c}
\]
\end{enumerate}
\end{rmk}

    Moreover, by \cite[Proposition 15.7]{MVW}, $\P^1$ is a symmetric object, so we can construct the stable $\infty$-category of tame motives (resp. tame motivic spectra) as\[
    \DM[t](k):=\Spt_{\P^1}(\DMeff[t](k))\quad \SH_t(k):=\Spt_{\P^1}(\SH^{S^1}_t(k))
    \]
using the general machinery of \cite{robalo} and \cite{HoveySpectra} (see also \cite[Section 1]{AnnalaIwasaUnivers}). In particular, we have the following result, whose proof is completely formal (see \emph{e.g.} \cite[\S 6.23]{CiDeg-LocalStableHomAlg}):
\begin{thm}\label{thm:ring-spectra}
Let $E_*$ be a graded commutative monoid in $\SH^{S^1}_t(k)$, together with a section $c\colon \cH^t(\P^1)\to E_1[2]$ such that for all $X \in \Sm_k$ and all $i$, the following composition is an equivalence:
\[
        \begin{tikzcd}
        \Map(\cH^t(X),E_i)\ar[r]\ar[ddrr,bend right=8] &\Map(\cH^t(\P^1_X),E_i\otimes \cH^t(\P^1)) \ar[r,"c"] &\Map(\cH^t(\P^1_X),E_i\otimes E_{1}[2])\ar[d,"\mu_{i,1}"]\\
        &&\Map(\cH^t(\P^1_X),E_{i+1}[2])\ar[d]\\
        &&\Map(\cH^t(\P^1_X,i_0),E_{i+1}[2]).
        \end{tikzcd}
\]
Then there is $\mathbf{E}\in \CAlg(\SH_t(k))$ such that for all $X\in \Sm_k$\[
\Map_{\SH_t(k)}(\Sigma^{\infty}(\cH^t(X)),\Sigma^{m,n}\mathbf{E}) \simeq R\Gamma_t(X,E_n[m]).
\]
Similarly, if $E_*$ is a graded commutative monoid in $\DMeff[t](k)$ as above, then there is a ring spectrum $\mathbf{E}$ in $\DM[t](k)$ representing $E_*$.
\end{thm}
By construction, if $\mathbf{E}\in \CAlg(\DM[t](k))$ representing $E_*$, then for all $q\geq 0$ the properties listed above imply:
\begin{enumerate}
    \item If $\cE \to X$ is a vector bundle of rank $r+1$, we have\[
        H^q_t(\cE,E_n)\xrightarrow{\simeq} H^q_t(X,E_n).
    \]
    and\[
    \bigoplus_{i=0}^r H^{q-2i}_t(X,E_{n-i})\cong H^q_t(\P(\cE),E_n).
    \]
    \item Let $X$ be a smooth scheme over $k$ and $Z$ a smooth closed
subscheme of $X$ of codimension $c$. Then the Gysin sequence gives\[
H^q_{t,Z}(X,E_n)\cong H^{q-2c}_{t}(Z,E_{n-c}).
\]
\end{enumerate}
We now compare tame with \'etale motives with invertible coefficients. First we record the following result, which is essentially \cite[Proposition 8.1]{HS2021}.
\begin{prop}\label{prop:tame-etale-transfers}
    Let $S$ be an $\F_p$-scheme, $X$ an $S$-scheme and $F$ be an \'etale sheaf of $\Z[1/p]$-modules on $X$. Then $H^q_{\et}(X,F)=H^q_t(X/S,F)$.  
\end{prop}
\begin{proof}
  The same proof of \emph{loc.cit.} works here: indeed if $R$ is henselian with residue field $k(R)$, then for all \'etale sheaves $G$ (not necessarily torsion) on $\Spec(R)$ we have $H^q_{\et}(\Spec(R),G) = H^q_{\et}(\Spec(k(R)),G_{k(R)})$: this is well known, see \cite[VII, Cor 8.6]{SGA4}. Then this allows to conclude that for every tame point $(\ol{x}_L,\ol{v}_L)$ of $X$ as in \emph{loc.cit.}\[
    H^q_{\et}(X_{(\ol{x}_L,\ol{v}_L)}^{th},F)\cong  H^q_{\et}(\Spec(L),F_{L}) = 0,
    \]
    since the Galois group of $L$ is a pro-$p$-group.
\end{proof}
\begin{lemma}\label{lem:tame-etale-sheafification}
Let $F$ be a tame sheaf of $\Z[1/p]$-modules with transfers. Then $F$
is also an \'etale sheaf with transfers.
    \begin{proof}
        The proof is completely analogous to \cite[Lemma 14.21]{MVW}, we report it for completeness. As in \emph{loc.cit.} we need to show that the kernel and cokernel of $F\to F_{\et}$ vanish, hence we may suppose that $F_{\et}=0$. If $F\neq 0$, there is $S=\Spec(R)\in \widetilde{\Sm_k}$ with $R$ tamely henselian and a nonzero element $c\in F(S)$. Since $F_{\et}=0$, there is a finite \'etale map $S'\to S$ with $c_{S'}=0$. Since $S$ is tamely henselian, the degree of $S'\to S$ is $p^r$ for $r\geq 0$ as the absolute Galois group of $k(S)$ is a pro-$p$-group, so since the map\[
        F(S)\to F(S')\to F(S)
        \]
        is the multiplication by $p^r$, which is invertible in $F(S)$, we conclude that $c=0$. This contradiction shows that $F=0$. 
    \end{proof}
\end{lemma}
\begin{cor}
    The sheafification map $\cD(\Sh_{t}^{tr}(k,\Z[1/p])\to \cD(\Sh_{\et}^{tr}(k,\Z[1/p])$ is an equivalence, therefore $\DMeff[t](k,\Z[1/p])\simeq \DMeff[\et](k,\Z[1/p])$ and $\DM[t](k,\Z[1/p])\simeq \DM[\et](k,\Z[1/p])$.
    \begin{proof}
        This is analogous to \cite[Proposition 2.2.9]{CD-etale} and follows from Proposition \ref{prop:tame-etale-transfers} and Lemma \ref{lem:tame-etale-sheafification}.
    \end{proof}
\end{cor}

Recall now the effective log motives $\logDMeff(k)$ and $S^1$-spectra $\logSH^{S^1}(k)$ defined in \cite{BPO} and \cite{BPO-SH}, with the localizations\[
\begin{tikzcd}
\logDMeff[\dNis](k)\ar[r,"L_{(\bcube,\loget)}"]&\logDMeff[\loget](k)\\
\logSH^{S^1}_{\dNis}(k)\ar[r,"L_{(\bcube,\loget)}"]&\SH^{S^1}_{\loget}(k).\\
\end{tikzcd}
\] 
We let $\cH_{\log}^{\tau}(-)$ and $\cM_{\log}^{\tau}(-)$ be the Yoneda functors similary as before. There are adjoint functors (see \cite[Construction 4.0.8]{BPO-SH})\[
\begin{tikzcd}
    \SH^{S^1}(k) \ar[r,"\omega^*"] &\logSH^{S^1}(k)\ar[l,shift left = 2,"\omega_*"]\ar[l,"\omega_\sharp"', shift right = 3]
\end{tikzcd}
\]
and similarly for $\DM$. Recall the assumption:
	\begin{nota}\label{nota:RS}
		We say that $k$ satisfies resolutions of singularities if the following two properties are satisfied (see \cite[Definition 7.6.3]{BPO} and \cite[Main Theorem I and II]{Hironaka}):
		\begin{enumerate}
			\item[(RS1)] For any integral scheme $X$ of finite type over $k$, there is a proper birational morphism $Y \to X$ of schemes over $k$, which is an isomorphism on the smooth locus, such that $Y$ is smooth over $k$.
			\item[(RS2)] Let $f \colon Y \to X$ be a proper birational morphism of integral schemes over $k$ such that $X$ is smooth over $k$ and let $Z_1, \ldots , Z_r$ be smooth divisors forming a strict normal crossing divisor on $X$. Assume that
			\[f^{-1}(X-Z_1 \cup\ldots \cup Z_r)\to X-Z_1 \cup\ldots \cup Z_r\] is an isomorphism. Then there is a sequence of blow-ups\[
			X_n\xrightarrow{f_{n-1}} X_{n-1} \xrightarrow{f_{n-2}}\ldots \xrightarrow{f_{0}}X_0\simeq X
			\]
			along smooth centers $W_i \subseteq X_i$ such that
			\begin{enumerate}[label=\alph*.]
				\item the composition $X_n \to X$ factors through $f$,
				\item $W_i$ is contained in the preimage of $Z_1 \cup \ldots \cup Z_r$ in $X_i$,
				\item $W_i$ has strict normal crossing with the sum of the reduced strict transforms of \[
				Z_1,\ldots,Z_r,f_0^{-1}(W_0),\ldots,f_{i-1}^{-1}(W_{i-1})\]
				in $X_i$.
			\end{enumerate}
		\end{enumerate}
	\end{nota}
    \begin{rmk}\label{rmk:comp}
        If $k$ satisfies (RS1) and (RS2) above, then for every scheme $X\in \Sm_k$ there exists $\ol{X}\in \Sm_k$ with an open immersion $j\colon X\subseteq \ol{X}$ such that $X-j(X)$ is the support of a simple normal crossing divisor. This divisor supports then a DF log structure $\partial X$ in the usual way (see \cite[Definition 1.7.1]{ogu}). The resulting log scheme $(X,\partial X)$ is called a \emph{smooth log compactification}.
    \end{rmk}
By \cite[Theorem 4.4 (2)]{DoosungA1invlog} (resp.  \cite[Proposition 8.2.8]{BPO}), if $k$ satisfies (RS1) and (RS2) then for all $X\in \Sm_k$ with smooth log compactification $(\ol{X},\partial X)$ as above, the image of $\cH^{\Nis}(X)$ (resp. $\cM^{\Nis}(X)$) via the above functor is $\cH_{\log}^{\dNis}(\ol{X},\partial X)$ (resp. $\cM_{\log}^{\dNis}(\ol{X},\partial X)$), in particular the functor $\omega^*$ is monoidal. 
\begin{lemma}\label{lem:tame-descent}
Let $k$ be a perfect field of characteristic $p$ that satisfies resolutions of singularities as in \ref{nota:RS}. Let $U\to X$ be a tame cover in $\Sm_k$ and for all $m$ let $U^m$ denote the $m$-fold fiber product $U\times_X U\ldots \times_X U$. Then the colimit along the \v Cech nerve induces an equivalence:\[
L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(X)\simeq \colim_m L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(U^m)\quad \textrm{in }\logSH^{S^1}_{\loget}(k)
\]
and similarly for $\logDMeff[\loget]$.
\end{lemma}
\begin{proof}
    We only do the proof for $\SH$, the proof of $\DM$ is similar (in fact easier). If $U\to X$ is finite, then this is \cite[Lemma 5.4]{mericicrys}. In general, we proceed by induction on the dimension of $X$. If the dimension is zero, then $U\to X$ is finite and we conclude. Since $U$ and $X$ are smooth separated of finite type over $\Spec(k)$, $U\to X$ is locally of finite presentation and separated. Since $U\to X$ is \'etale, it is flat and locally quasi-finite (see \cite[\href{https://stacks.math.columbia.edu/tag/03WS}{Tag 03WS}]{stacks-project}), and since $U$ is quasi-compact it has universally bounded fibres by \cite[\href{https://stacks.math.columbia.edu/tag/03JA}{Tag 03JA}]{stacks-project}, so by \cite[\href{https://stacks.math.columbia.edu/tag/07RY}{Tag 07RY}]{stacks-project} there is a closed subscheme $Z\subseteq X$ such that $U_{|X-Z}\to X-Z$ is finite (we consider $Z_{n-1}$ in \emph{loc.cit.}). Let $d$ be the relative codimension of $Z$ in $X$ and let $Z_0\subset Z$ be the singular locus of $Z$, so that $Z-Z_0\subseteq X-Z_0$ is a closed immersion in $\Sm_k$. Let $V\to X-Z_0$ be an trivializing open for the normal bundle of $Z-Z_0\subseteq X-Z_0$: by taking $V$ small enough, we can suppose that $(Z-Z_0)$ has pure codimension $d$ inside $V$.
    Let $V':= V\times_X(Z-Z_0)$ and $U' = U\times_X V\times_X(Z-Z_0)$. Then since $U^m\times_X V\to V$ and $V\to X$ are all flat, the normal bundle commute with the pullback so the Gysin sequences imply that the columns of the following diagram are fiber sequences:\[
    \begin{tikzcd}
\colim_m L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((U^m)_{|V-Z_{|V}})\ar[r]\ar[d] &L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(V-Z_{|V})\ar[d]\\
        \colim_m L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((U^m)_{|V})\ar[r]\ar[d] &L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(V)\ar[d]\\
        (\colim_m L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((\P^{d}/\P^{d-1})\times (U')^m)\ar[r]&L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(\P^{d}/\P^{d-1})\times V').
    \end{tikzcd}
    \]
Since both $\omega^*$ and $L_{(\bcube,\loget)}$ are monoidal, we can take $\P^{d}/\P^{d-1}$ outside and by induction on dimension the bottom horizontal map is an equivalence, and the top horizontal map is an equivalence again by \cite[Lemma 6.4]{mericicrys} since $U_{|V-Z_{|V}}\to V-Z_{|V}$ is finite, so the middle horizontal map is an equivalence. We can then take $V\to X$ a trivializing cover and apply the diagram above to $V^r:= V\times_X V\times_X\ldots \times_X V$, so by Nisnevich descent of $\omega^*$ (see \cite[Proposition 3.7]{DoosungA1invlog}) we deduce that\[
\begin{aligned}
&\colim_{m} L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((U^m)_{|X-Z_0})\simeq \colim_r \colim_{m} L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((U ^m)_{|V^r})\\
&\simeq \colim_r L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(V^r) \simeq L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(X-Z_0).
\end{aligned}
\]  
By taking now $Z_n$ as the singular locus of $Z_{n-1}$, we get a chain of strict closed subschemes such that \[
\colim L_{(\bcube,\loget)}\omega^*\cH^{\Nis}((U^\bullet)_{|X-Z_n})\simeq L_{(\bcube,\loget)}\omega^*\cH^{\Nis}(X-Z_n)\textrm{ for all }n,
\] 
so we conclude since $X$ has finite Krull dimension. 
\end{proof}
For the rest of the section, we assume that $k$ satisfies resolutions of singularities as in \ref{nota:RS}. We are now ready to prove our main theorem:
\begin{thm}\label{thm:comparison}
    The adjunction\[
    \begin{tikzcd}
    \SH^{S^1}(k) \ar[r,"\omega^*", shift left = 2] &\logSH^{S^1}(k)\ar[l,"\omega_*"]
    \end{tikzcd}
    \]
    induces an adjunction\[
    \begin{tikzcd}
    \SH^{S^1}_t(k) \ar[r,"\omega^*_t", shift left = 2] &\logSH^{S^1}_{\loget}\ar[l,"\omega_*^t"](k)
    \end{tikzcd}
    \]
fitting in commutative diagrams:\[
    \begin{tikzcd}
    \SH^{S^1}(k) \ar[r,"\omega^*"]\ar[d,"L_{(\A^1,t)}"] &\logSH^{S^1}(k)\ar[d,"L_{(\bcube,\loget)}"] && 
    \SH^{S^1}(k) &\logSH^{S^1}(k)\ar[l,"\omega_*"]\\
     \SH^{S^1}_t(k) \ar[r,"\omega^*_t"] &\logSH^{S^1}_{\loget}(k) &&
    \SH^{S^1}_t(k) \ar[u,"i_{(\A^1,t)}"]&\logSH^{S^1}_{\loget}\ar[l,"\omega_*^t"](k)\ar[u,"i_{(\bcube,\loget)}"]
    \end{tikzcd}
\]
and similarly for $\DMeff$. The functor $\omega^*_t$ is monoidal.
\begin{proof}
    The functor $L_{(\bcube,\loget)}\omega^*$ satisfies $\A^1$-invariance by construction and tame descent by Theorem \ref{lem:tame-descent} above, therefore it factors through $L_{(\A^1,t)}$ inducing $\omega^*_t$ that fits in the first commutative square. Since $L_{(\bcube,\loget)}\omega^*$ is a left adjoint, it preserves all colimits. Since both $i_{(\A^1,t)}$ and $i_{(\bcube,\loget)}$ are fully faithful, we have that for $\{M_i\}_{i\in I}$ a system in $\DMeff[t](k,\Z)$:
    \begin{align*}
            &\colim_{i\in I} \omega^*_t M_i \simeq \colim_{i\in I} \omega^*_tL_{(\A^1,t))}i_{(\A^1,t)}M_i \simeq  
            \colim L_{(\bcube,\loget))}\omega^*i_{(\A^1,t)} M_i \simeq 
            L_{(\bcube,\loget))}\omega^* \colim i_{(\A^1,t)}M_i \\
            &\simeq \omega^*_t L_{(\A^1,t))}\colim i_{(\A^1,t)} M_i \simeq \omega^*_t \colim_{i\in I} M_i,
    \end{align*}
    where the last equivalence follows from the fact that $L_{(\A^1,t)}\dashv i_{(\A^1,t)}$ is a localization, so by the adjoint functor theorem there exists a right adjoint $\omega_*^t$, which fits by adjunction in the right square. The monoidality of $\omega^*_t$ then follows from the monoidality of $L_{(\bcube, \loget)}$ and $L_{(\A^1, t)}$, which holds by definition of the monoidal structure on tame motives, and the monoidality of $\omega^*$, which holds by \cite[Theorem 4.4(4)]{DoosungA1invlog}.
\end{proof}
\end{thm}
\begin{cor}
    The integral $p$-adic cohomology of \cite{mericicrys} factors through $\DMeff[t](k)$, inducing a realization\[
    R\Gamma_p^t\colon \DMeff[t](k)\to \cD(R(k))\quad X\mapsto \lim_m R\Gamma((\ol{X},\partial X), W_m\Lambda^\bullet),
    \]
    where $W_m\Lambda^\bullet$ is the logarithmic de Rham--Witt complex of Hyodo--Kato (see \cite{mokrane} or \cite{matsuue}) and $R(k)$ is the Raynaud ring of Ekhedal \cite{EkedahlII}.
    \begin{proof}
        $R\Gamma_p$ is defined as:\[
        \lim_m \Map_{\logDMeff[\loget](k,\Z)}(L_{(\bcube,\loget)}\omega^*(-),W_m\Lambda^\bullet)\colon \DMeff(k)\to \cD(R(k))
        \]
        and so by Theorem \ref{thm:comparison} this agrees with\[
        \DMeff(k)\xrightarrow{L_{(\A^1,t)}}  \DMeff[t](k)\xrightarrow{\lim_m R\Gamma(\omega^*_t(-), W_m\Lambda^\bullet)} \cD(R(k)).
        \]
    \end{proof}
\end{cor}

\begin{cor}
Let $F$ be a strictly $(\bcube,\loget)$-invariant sheaf of abelian groups, so that $F[0]\in \logSH^{S^1}_{\loget}$. Let $G\in \Sh_{t}(\Sm_k,\Spt)$ be the underlying sheaf of spectra of $\omega_*^t F[0]$. Then the tame sheaf $\pi_0G$ is strongly $\A^1$-invariant, \emph{i.e.} for all $X\in \Sm_k$\[
\pi_0G(\A^1_X)\simeq \pi_0G(X)\textrm{ and }H^1_t(\A^1_X,\pi_0G)\simeq H^1_t(X,\pi_0G).
\] 
\begin{proof}
    Since $\omega_*$ is left $t$-exact with respect to the homotopy $t$ structures of $\logSH^{S^1}(k)$ and $\SH^{S^1}(k)$ and the inclusion $R\epsilon_*\colon\Sh_t(\Sm_k,\Spt)\to \Sh_{\Nis}(\Sm_k,\Spt)$ is fully faithful and left $t$-exact for the usual $t$-structure induced by the Postnikov $t$-structure on $\Spt$, we deduce that $\tau_{\geq 1}G = 0$, so we have a fiber sequence in $\Spt$:\[
    \pi_0G(Y)\to G(Y)\to \tau_{< 0}G(Y),
    \] 
    which implies that $\pi_0G(Y) = \pi_0(G(Y))$. Since $G$ is $(\A^1,t)$-local, we conclude that $\pi_0(G(\A^1_X)) = \pi_0(G(X))$. Finally, consider the hypercohomology speactral sequence, which is functorial in $Y$:\[
    H^p_t(Y,\pi_{-q}G)\Rightarrow H^{p+q}_t(X,G)
    \]
    whose five term exact sequence gives an injective map\[
    H^1_t(Y,\pi_0G)\hookrightarrow H^1_t(Y,G).
    \]
    Since $G$ is $(\A^1,t)$-local, we have that $H^1_t(\A^1_X,G) = H^1_t(X,G)$, so the map $H^1_t(\A^1_X,\pi_0G) \to H^1_t(X,\pi_0G)$ induced by the zero section is injecitve. On the other hand, this map has a retraction induced by the projection $\A^1_X\to X$, so it is also surjective.
\end{proof}
\end{cor}

\begin{example}\label{ex:drW}
    Consider the \'etale sheaves $\nu_m(n)$: they fit in an left exact sequence in $\mathbf{RSC}_{\Nis}$:\[
    0\to \nu_m(n)\to W_m\Omega^n\xrightarrow{F-1}W_m\Lambda^n/dV^{m-1}\Omega^{n-1}
    \] 
    so by \cite{shujilog} and \cite[Theorem 4.4]{mericicrys} they fit in a left exact sequence of log \'etale sheaves:\[
        0\to \widetilde{\nu_m(n)}\to W_m\Lambda^n \xrightarrow{F-1}W_m\Lambda^n/dV^{m-1}\Lambda^{n-1}
    \]
    where $\widetilde{\nu_m(n)}$ is as in \eqref{eq:nu-tilde}.
    By \cite[Proposition 2.13]{lorenzon}, the last map is surjective in the log \'etale topology. Moreover, $W_m\Lambda^n$ and $W_m\Lambda^n/dV^{m-1}\Lambda^{n-1}$ are both  strictly $(\bcube,\loget)$-invariant by \cite[Theorem 4.2]{mericicrys}, hence $\widetilde{\nu_m(n)}$ are strictly $(\bcube,\loget)$-invariant, which implies that $L_{\loget}\omega^*\nu_m(n)\simeq \widetilde{\nu_m(n)}[0]$ is $(\bcube,\loget)$-local. This implies that $\pi_0\omega_*^tL_{\loget}\omega^*\nu_m(n)$ is strongly $(\A^1,t)$-invariant, but for all $X\in \Sm_k$ with smooth log compactification $(\ol{X},\partial X)$ as in Remark \ref{rmk:comp} we have that:
    \[
        \pi_0\omega_*^tL_{\loget}\omega^*\nu_m(n)(X) = \widetilde{\nu_m(n)}(X,\partial X) = \nu_m(n)(X),
    \]
    so we deduce that the sheaves $\nu_m(n)$ are strongly $(\A^1,t)$-invariant.
\end{example}

\section{\texorpdfstring{$\A^1$}{A1}-invariance of higher tame cohomology}

In this section, we use the result of \cite{amine} to prove the $\A^1$-invariance of higher tame cohomology. Notice that the assumption on resolutions of singularities on $k$ are still needed. 

\begin{lemma}\label{lem:pbf}
For all $m,n$, there is an equivalence in $\cD(\Sh_t^{\rm tr}(\Sm_k,\Z))$:\[
\nu_m(n)[0]\oplus \nu_m(n-1)[-1]\simeq R\uHom_{\Sh_{t}^{\tr}}(\Z_{\tr}(\P^1),\nu_m(n)).\] 
\end{lemma}
\begin{proof}
    Recall that $\nu_m(n)$ are $\A^1$-invariant Nisnevich sheaves with transfers and $(\nu_m(n))_{-1}\cong \nu_m(n-1)$, so the projective bundle formula in $\DMeff(k,\Z)$ (see \cite[Theorem 15.1 and Proposition 24.8]{MVW}) gives by adjunction maps in $\cD(\Sh_{\Nis}^{\rm tr}(\Sm_k,\Z))$\[
    (\nu_m(n)[0]\oplus \nu_m(n-1)[-1])\otimes \Z_{\rm tr}(\P^1) \to \nu_m(n).
    \]
    Since the tame sheafification is monoidal, again by adjunction this gives a map in $\cD(\Sh_{t}^{\rm tr}(k,\Z))$\[
     \nu_m(n)[0]\oplus \nu_m(n-1)[-1] \to R\uHom_{\Sh_t^{\rm tr}}(\Z_{\tr}(\P^1),\nu_m(n)).
    \]
    Let $X\in \Sm_k$: we show that the induced map\[
    R\Gamma(X,\nu_m(n))\oplus R\Gamma(X,\nu_m(n-1))[-1]\to R\Gamma(\P^1_X,\nu_m(n))
    \] 
    is an equivalence. Let $X\in \Sm_k$ and let $\ol{X}$ smooth and proper with $j\colon X\hookrightarrow \ol{X}$ open and $\ol{X}-j(X)$ supported on a simple normal crossing divisor $D=D_1+\ldots D_t$. Then we proceed by double induction on $\dim(X)$ and $t$. If $\dim(X)=0$ or $t=0$, then $X$ is proper so the map above is an equivalence by \cite[Proposition 8.2]{HS2021} and the projective bundle formula for the \'etale cohomology of $\nu_m(n)$ by \cite{Gros1985}. In general, let $X':= \ol{X}-(D_1\cup \ldots D_{t-1})$ and $D':= D_t-(D_1\cup \ldots D_{t-1})$: by \cite[Theorem 1.3.1.]{amine} we have a map of long exact sequences\[
    \begin{tikzcd}
    H^{q-2}_t(D',\nu_m(n-2))\oplus H^{q-1}_t(D',\nu_m(n-1))\ar[r,"(1)"]\ar[d]&H^{q-1}_t(\P^1_{D'},\nu_m(n-1))\ar[d]\\
        H^{q-1}_t(X',\nu_m(n-1))\oplus H^{q}_t(X',\nu_m(n)) \ar[r,"(2)"]\ar[d]&H^q_t(\P^1_{X'},\nu_m(n))\ar[d]\\
        H^{q-1}_t(X,\nu_m(n-1))\oplus H^q_t(X,\nu_m(n))\ar[r,"(3)"]\ar[d]&H^q_t(\P^1_X,\nu_m(n))\ar[d]\\
        H^{q-1}_t(D',\nu_m(n-2))\oplus H^{q}_t(D',\nu_m(n-1))\ar[r,"(4)"]\ar[d]&H^{q}_t(\P^1_{D'},\nu_m(n-1))\ar[d]\\
        H^{q}_t(X',\nu_m(n-1))\oplus H^{q+1}_t(X',\nu_m(n)) \ar[r,"(5)"]&H^{q+1}_t(\P^1_{X'},\nu_m(n))
    \end{tikzcd}
    \]
    By induction hypotheses the maps $(1),(2),(4),(5)$ are isomorphisms, so $(3)$ is also an isomorphism.
\end{proof}
\begin{cor}\label{cor:pbf-tamely-hens}
    Let $X$ be the tame henselization of a smooth scheme over $k$ at a tame point $(\ol{x},v)$. Then $H^q_t(\P^1_X,\nu_m(n))=0$ for $q\geq 2$. 
    \begin{proof}
        We have $H^q_t(\P^1_X,\nu_m(n))\cong H^q_t(X,\nu_m(n))\oplus H^{q-1}_t(X,\nu_m(n-1))$ by Lemma \ref{lem:pbf} and \cite[Theorem 4.7]{HS2021}, and the right hand side is zero for $q\geq 2$ since $X$ is tamely henselian.
    \end{proof}
\end{cor}
\begin{thm}\label{thm:A1-inv-higher-cohomology}
    Let $X$ be the tame henselization of a smooth scheme over $k$ at a tame point $(\ol{x},v)$. Let $U\subseteq \P^1_{\ol{k}}$ be the complement of finitely many closed points. Then $H^q_t(U_X,\nu_m(n))=0$ for $q\geq 2$. 
    \begin{proof}
        Let $U=\P^1_{\ol{k}}-\{x_1\ldots x_n\}$. Since $X=\Spec(A)$ is tamely henselian with respect to $v$ trivial on $k$, we have that $\ol{k}\hookrightarrow A$, therefore $U_X \cong \P^1_{\ol{k}}-\{X_1\ldots X_n\}$, where $X_i=X\times_{\ol{k}}{x_i}$ is isomorphic to $X$, hence it is tamely henselian. Since $X$ is a cofiltered limit of smooth schemes with affine transition morphisms, \cite[Theorem 1.3.1]{amine} and \cite[Theorem 4.6]{HS2021} give a long exact sequence\[
        \ldots \to H^q_t(\P^1_X,\nu_m(n))\to H^q_t(U_X,\nu_m(n))\to \oplus H^{q}(X_i,\nu_m(n-1))\to \ldots
        \]
        Then $\oplus H^{q}_t(X_i,\nu_m(n-1))=0$ since each $X_i$ is tamely henselian and $H^q_t(\P^1_X,\nu_m(n))=0$ by Corollary \ref{cor:pbf-tamely-hens}, so the theorem follows.
    \end{proof}
\end{thm}

\section{Proof of the main theorems}

In this section, we put together the results and prove the theorems in the introduction. 
Let $k$ be a field that satisfies resolutions of singularities as in \ref{nota:RS}. Then combining Example \ref{ex:drW} and Theorem \ref{thm:A1-inv-higher-cohomology}, we conclude that for all $X$ tame henselization of a smooth scheme at a tame point $(\ol{x},v)$ and all $q\geq 0$ we have
\begin{equation}\label{eq:A1-inv}
    H^q_t(\A^1_X,\nu_m(n))\simeq H^q_t(X,\nu_m(n))
\end{equation}
We are now ready to prove Theorem \ref{thm:motivic-dRW}, \emph{i.e.} the existence of the motivic ring spectrum $H\Z/p^m$. By the usual spectral sequence argument \eqref{eq:A1-inv} implies that $\nu_m(n)$ is a strictly $\A^1$-invatiant tame sheaf with transfers, in particular the collection $\nu_m(*)[-*]$ gives rise to a graded commutative monoid in $\DMeff[t](k)$. Then to conclude it is enough to show that the maps induced by the Chern classes\[
R\Gamma_t(X,\nu_m(n))\oplus R\Gamma_t(X,\nu_m(n-1))[1]\to R\Gamma_t(\P^1_X,\nu_m(n))
\]
are equivalences, which follows from Lemma \ref{lem:pbf}. Then the spectrum $H\Z/p^m$ exists by Theorem \ref{thm:ring-spectra}
\begin{rmk}\label{rmk:GL}
In fact, by Geisser--Levine \cite[Theorem 1.1]{GL}, there is a quasi-isomorphism of complexes of Nisnevich sheaves with transfers: $\Z/p^m(n)\simeq \nu_m(n)[-n]$, where the left hand side is the motivic complex. This implies that the graded $E_{\infty}$-ring in $\DMeff[\Nis](k,\Z/p^m)$ given by the collection $\nu_m(*)[-*]$ builds up the unit of $\DM[\Nis](k,\Z/p^m)$. Since the localization functor $L_{(\A^1,t)}^{st}\colon \DM[\Nis](k,\Z/p^m)\to \DM[t](k,\Z/p^m)$ is monoidal, it preserves the unit: this implies that the ring spectrum $H\Z/p^m$ is the unit of $\DM[t](k,\Z/p^m)$, since for all $m,n$ we have that $L_{(\A^1,t)}\nu_m(n) = L_{t}\nu_m(n)$.
\end{rmk}
As observed in the introduction, Theorem \ref{thm:main-intro} is now a mere consequence of Theorem \ref{thm:mot-intro} and the motivic properties of $\DM$ (see \eqref{item:1} and \eqref{item:2}).
Finally, we prove Theorem \ref{thm:comparison}, which we restate for the sake of the reader:
\begin{thm}
Let $k$ be a perfect field of characteristic $p$ that satisfies resolutions of singularities as in \ref{nota:RS}. For all $X\in \Sm_k$ with smooth log compactification $(\ol{X},\partial X)$ as in Remark \ref{rmk:comp}, we have that\[
R\Gamma_t(X,\nu_m(n))\simeq R\Gamma_{\loget}((\ol{X},\partial X),\widetilde{\nu_m(n)})
\]
\begin{proof}    
The canonical map $\nu_m(n)[0]\to \omega_*^t\omega^*_t\nu_m(n)[0]$ induces a map
\begin{align*}
R\Gamma_t(X,\nu_m(n))\to &\Map_{\DM[t]}(\cM^t(X), \omega_*^t\omega^*_t\nu_m(n)[0])\simeq \\
&\Map_{\logDM[\loget]}(\cM^{\loget}(\ol{X},\partial X),L_{(\bcube,\loget)}\omega^*\nu_m(n)[0])
\end{align*}
functorial in $X$ and $(\ol{X},\partial X)$, and since $L_{\loget}\omega^*\nu_m(n)[0]\simeq \widetilde{\nu_m(n)}[0]$ is already $(\bcube,\loget)$-local, we have that $L_{(\bcube,\loget)}\omega^*\nu_m(n)[0]\simeq \widetilde{\nu_m(n)}[0]$, therefore the map above induces a map \[R\Gamma_t(X,\nu_m(n))\to R\Gamma_{\loget}((\ol{X},\partial X),\widetilde{\nu_m(n)}).\] We will show that this map is an equivalence. 
Let $|\partial X| = D_1 + \ldots D_r$: we proceed by double induction on $\dim(X)$ and $r$. If $r=0$, then $X=\ol{X}$ is proper, so\[
H^q_t(X,\nu_m(n))\simeq H^q_{\et}(X,\nu_m(n)) = H^q_{\loget}((X,\triv),\widetilde{\nu_m(n)}).
\]
If $\dim(X)=0$, then $\partial X=0$ so it follows from the case above. For $\dim(X)>0$ and $r(X)>0$, let $\partial X'$ be the log structure supported on $D_1+\ldots D_{r-1}$ and $X':=\ol{X}-|\partial X'|$: then Morel--Voevodsky purity of \ref{item:2} in \ref{rmk:pbf-gysin} gives the fiber sequence \[
\cM^t(X)\to \cM(X')\to \cM^t(D_1\cap X')\otimes \cM^t(\P^1,i_0),
\]
so using the computation of the cohomology of $\P^1$ in Lemma \ref{lem:pbf} and \cite[Corollary 4.5]{mericicrys} gives for all $q$ a diagram whose columns are long exact sequences:\[
\begin{tikzcd}
    H^{q-1}_t(D_1\cap X',\nu_m(n-1))\ar[r,"(1)"]\ar[d]& H^{q-1}_{\loget}((D_1\partial X_{D_1}),\widetilde{\nu_m(n-1)})\ar[d]\\
    H^q_t(X',\nu_m(n))\ar[r,"(2)"]\ar[d]&H^q_{\loget}((X,\partial X'),\widetilde{\nu_m(n)})\ar[d] \\
    H^q_t(X,\nu_m(n))\ar[r,"(3)"]\ar[d]& H^q_{\loget}((X,\partial X),\widetilde{\nu_m(n)})\ar[d]\\
    H^q_t(D_1\cap X',\nu_m(n-1))\ar[r,"(4)"]\ar[d]& H^q_{\loget}((D_1\partial X_{D_1}),\widetilde{\nu_m(n-1)})\ar[d]\\
    H^{q+1}_t(X',\nu_m(n))\ar[r,"(5)"] &H^{q+1}_{\loget}((X,\partial X'),\widetilde{\nu_m(n)})\\
    &&&&
\end{tikzcd}
\]
By induction on the dimension, $(1)$ and $(4)$ are isomorphisms, and by induction on $r$ $(2)$ and $(5)$ are isomorphisms, so $(3)$ is an isomorphism, concluding the proof.
\end{proof}
\end{thm}

\bibliographystyle{alpha}
\bibliography{bibMerici}

\end{document}